PLEASE ANSWER IN 30 MINUTES

PM.53 Sun Oracle is launching a new computer component. The team has broken the project down into activities, with the following durations:

Activity     Duration (weeks)     Immediate Predecessor(s)  
A 20 -
B 12 -
C 15 A
D 5 B
E 6 D
F 14 A,E
G 5 F
H 3 C,G


What is the duration of the project?
   weeks

What is the latest start week for activity C?
   

What is the slack time for activity C?
   weeks

If activity D were shortened by 3 weeks, what would be the new duration of the project?
   weeks
Answer #1
Activity ES EF LS LF Slack
A 0 20 3 23 3
B 0 12 0 12 0
C 20 35 27 42 7
D 12 17 12 17 0
E 17 23 17 23 0
F 23 37 23 37 0
G 37 42 37 42 0
H 42 45 42 45 0

Slack = LF-EF = LS-ES

Duration of the project = duration of the longest path B-D-E-F-G-H = 45 weeks

latest start week for activity C = 27

Slack time for activity C = 7 weeks

If activity D were shirtened by 3 weeks new duration of the project would be 45-3 = 42 weeks as activity D is in critical path